OpenStudy (anonymous):

Simplify. Give answers using positive exponents. Leave your answer in exponential notation. (3^6)^-9

OpenStudy (jhannybean):

let's say we take our (3^6) and substitute that for "f" another way to write f ^-9 is by putting it under 1. 1/f^9

OpenStudy (anonymous):

3^-3 maybe

OpenStudy (jhannybean):

now that it's rewritten as 1/f^9 , we can re substitute " f " back in, so we would have 1/(3^6)^9 = 1/3^(6 * 9) can you solve this now? :)

OpenStudy (jhannybean):

remember: \(\large (a^x)^y = a^{x*y}\)

OpenStudy (jhannybean):

where x and y are exponents.

OpenStudy (anonymous):

Ok

OpenStudy (jhannybean):

what did you get for your answer?

OpenStudy (anonymous):

1/3^-54

OpenStudy (jhannybean):

remember how we took (3^6) as f? and then we had ( f )^-9? in order to make ( f )^-9 positive, we bring f AND the exponent under a 1, to make it a fraction. so therefore we would have 1/ ( f )^9 Are you following?

OpenStudy (anonymous):

ok 1/3^54 thank-you kindly

OpenStudy (jhannybean):

no problem :) do you understand how this works now? :)

OpenStudy (anonymous):

Indeed I do. Many thanks again

OpenStudy (jhannybean):

Great! :D